# DiffSampling: Enhancing Diversity and Accuracy in Neural Text Generation

### v4

@@ D.4 Ablation Study on the Dynamic Upper Bound
Appendix E Qualitative Analysis
E.1 Divergent Association Task
@@
E.2 Math Problem Solving
Tables 21 and 22 report two qualitative examples of our DiffSampling methods for the GSM8K test set (preferred over MATH due to output length). The first thing we can notice is how a temperature of $10.0$ (and occasionally a temperature of $2.0$) makes the baselines generate random tokens, while our methods remain always on topic (even though potentially varying in the final result). In particular, temperature scaling on DiffSampling-cut has the effect of rephrasing some sentences, but never losing the overall meaning and mathematical steps.
E.3 Extreme Summarization
Instructed Model. Tables 23 and 24 report some qualitative examples of our DiffSampling methods for XSum when adopting the instructed model. Again, higher temperatures make top-$p$ and min-$p$ behave more randomly. Our methods show less variety and often produce a similar output, but remain consistent across all tested temperatures.
Pre-Trained Model. Tables 25 and 26 report some qualitative examples of our DiffSampling methods for the XSum dataset when adopting the pre-trained model. Similar to what was experienced for the instructed model, top-$p$ and min-$p$ fail in producing coherent and meaningful outputs at higher temperatures, and sometimes they fail even at a temperature of $1.5$. While the pre-trained model is more prone to less coherence, our methods usually generate appropriate summaries, and on the rare occasions they fail to do so, the output is still somehow connected to the input text or the request.
E.4 Story Generation
Instructed Model. Tables 27 and 28 report some qualitative examples for Llama-3.2-3B-Instruct. As apparent, our two relaxations sometimes behave very close to top-$p$ and min-$p$, while in other cases they start diverging quite soon, while preserving the general meaning of the output. This confirms, once again, how our methods perform subtle corrections, extending the range of possible tokens occasionally but meaningfully. Instead, DiffSampling-cut diverges almost immediately from the greedy strategy, as we would expect for a creative task where multiple outcomes are equally acceptable.
Pre-Trained Model. Finally, Tables 29 and 30 report some qualitative examples for Llama-3.2-3B. As guessed from the quantitative scores, the greedy strategy tends to repeat the same tokens, with very poor variability. Also DiffSampling-cut tends to repeat sentences multiple times, but this happens less frequently, and sometimes they are broken by different tokens, most likely due to our cutting strategy (see how the same sentence contains any time a different subject in Table 30). Our other two methods are less prone to repetition and write more coherent text, while diverging sooner from their most similar baselines.

### v5

@@ G.2 Ablation Study on the Lower Bound
To sum up, when greediness is desirable, a lower value of $p_{lb}$ can lead to high quality and diversity; otherwise, increasing $p_{lb}$ improves diversity, but the cost in terms of validity is not negligible and requires careful consideration. We suggest practitioners select the most appropriate $p_{lb}$ value by running it on a validation set if available, and otherwise lie in the $[0.8,0.95]$ range, which has shown competitive results on both quality and diversity metrics.
G.3 Ablation Study on the Dynamic Upper Bound
@@
Table 22 reports the results for the math problem-solving tasks, considering the GSM8K (left side) and MATH (right side) test sets. As expected, the against-greedy diversity scores and cross-input EAD decrease together with $p_{min}$, plateauing at $p_{min}=0.6$ (from that on, results are comparable with DiffSampling-cut); specularly, accuracy is lower at smaller $p_{min}$, but the instructed model reaches a competitive score even at $p_{min}=0.3$.
@@
The same holds for XSum as well. As shown in Table 23, diversity decreases when increasing $p_{min}$ and plateaus at 0.5, while quality rapidly increases for the pre-trained model and is almost constant for the instructed model.
@@
To sum up, values above $0.5$ are not different from DiffSampling-cut, while lower $p_{min}$ can help foster diversity with a small loss in accuracy, especially for instructed models. We suggest practitioners select the most appropriate $p_{min}$ value by running it on a validation set if available, and otherwise lie in the $[0.05,0.3]$ range, with a lower or higher value depending on whether it is preferable to have more diversity or quality, respectively.
Appendix H Qualitative Analysis
@@
H.1 Divergent Association Task
@@
H.2 Math Problem Solving
Tables 25 and 26 report two qualitative examples of our DiffSampling methods for the GSM8K test set (preferred over MATH due to output length). The first thing we can notice is how a temperature of $10.0$ (and occasionally a temperature of $2.0$) makes the baselines generate random tokens, while our methods remain always on topic (even though potentially varying in the final result). In particular, temperature scaling on DiffSampling-cut has the effect of rephrasing some sentences, but never losing the overall meaning and mathematical steps.
H.3 Extreme Summarization
Instructed Model. Tables 27 and 28 report some qualitative examples of our DiffSampling methods for XSum when adopting the instructed model. Again, higher temperatures make top-$p$ and min-$p$ behave more randomly. Our methods show less variety and often produce a similar output, but remain consistent across all tested temperatures.
Pre-Trained Model. Tables 29 and 30 report some qualitative examples of our DiffSampling methods for the XSum dataset when adopting the pre-trained model. Similar to what was experienced for the instructed model, top-$p$ and min-$p$ fail in producing coherent and meaningful outputs at higher temperatures, and sometimes they fail even at a temperature of $1.5$. While the pre-trained model is more prone to less coherence, our methods usually generate appropriate summaries, and on the rare occasions they fail to do so, the output is still somehow connected to the input text or the request.
H.4 Story Generation
Instructed Model. Tables 31 and 32 report some qualitative examples for Llama-3.2-3B-Instruct. As apparent, our two relaxations sometimes behave very close to top-$p$ and min-$p$, while in other cases they start diverging quite soon, while preserving the general meaning of the output. This confirms, once again, how our methods perform subtle corrections, extending the range of possible tokens occasionally but meaningfully. Instead, DiffSampling-cut diverges almost immediately from the greedy strategy, as we would expect for a creative task where multiple outcomes are equally acceptable.
Pre-Trained Model. Finally, Tables 33 and 34 report some qualitative examples for Llama-3.2-3B. As guessed from the quantitative scores, the greedy strategy tends to repeat the same tokens, with very poor variability. Also DiffSampling-cut tends to repeat sentences multiple times, but this happens less frequently, and sometimes they are broken by different tokens, most likely due to our cutting strategy (see how the same sentence contains any time a different subject in Table 34). Our other two methods are less prone to repetition and write more coherent text, while diverging sooner from their most similar baselines.